There are three numbers in the ratio of 2:3:4. The sum of their squares is 464. Find the numbers.
Answer by rfer(16322)

2x^2+3x^2+4x^2=464
9X^2=464
x^2=51.56
x=7.18
--------------
2x^2=103.1
3x^2=154.7
4x^2=206.2
total=464